OpenStudy (anonymous):

how do i solve this? 2(3x - 4)

OpenStudy (anonymous):

multiply each term by 2 i.e (2*3x)-(4*2) =6x-8

OpenStudy (mathstudent55):

The term "solving" applies to equations. For that you need an equal sign. If all you're given is what you wrote above, then perhaps you need to simplify it. Then, look above at moli's answer

OpenStudy (anonymous):

The property by which we do this is called DISTRIBUTIVE PROPERTY which says that: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

it makes sense now thanks